# Intelligence Foundation Model: A New Perspective to Approach Artificial General Intelligence

### v2

@@ Abstract
We propose a new perspective for approaching artificial general intelligence (AGI) through an intelligence foundation model (IFM). Unlike existing foundation models (FMs), which specialize in pattern learning within specific domains such as language, vision, or time series, IFM aims to acquire the underlying mechanisms of intelligence by learning directly from diverse intelligent behaviors. Vision, language, and other cognitive abilities are manifestations of intelligent behavior; learning from this broad range of behaviors enables the system to internalize the general principles of intelligence. Based on the fact that intelligent behaviors emerge from the collective dynamics of biological neural systems, IFM consists of two core components: a novel network architecture, termed the state neural network, which captures neuron-like dynamic processes, and a new learning objective, neuron output prediction, which trains the system to predict neuronal outputs from collective dynamics. The state neural network emulates the temporal dynamics of biological neurons, allowing the system to store, integrate, and process information over time, while the neuron output prediction objective provides a unified computational principle for learning these structural dynamics from intelligent behaviors. Together, these innovations establish a biologically grounded and computationally scalable foundation for building systems capable of generalization, reasoning, and adaptive learning across domains, representing a step toward truly AGI.
@@ I Introduction
Artificial general intelligence (AGI) refers to systems capable of understanding, learning, and applying knowledge across a wide range of tasks and domains at a level comparable to or exceeding that of humans [agi]. Despite decades of research, AGI has not yet been realized. To advance toward AGI, we propose a novel perspective through an Intelligent Foundation Model (IFM). The main idea of IFM is to learn “intelligence” from a broad range of intelligent behaviors, rather than specific manifestations such as language, vision, or other cognitive tasks. To date, the only known instance of such general intelligence is human intelligence [humanagi]. IFM therefore aims to capture human intelligence by learning from diverse human intelligent behaviors to approach AGI. This approach is grounded in the observation that every human intelligent behavior, whether involving perception, reasoning, or decision-making, can ultimately be represented as a temporal sequence of neuronal input–output transformations in human brain. By framing intelligence in terms of these observable sequences, IFM transforms the abstract and elusive goal of “learning intelligence” into a concrete sequence learning problem. Following the paradigm of FMs [fm], IFM utilizes large-scale neuronal input-output transformation data that comprehensively represent diverse human intelligent behaviors. From these data, we expect that IFM can learn the universal underlying mechanisms that give rise to human cognitive abilities, enabling artificial systems to generalize across tasks and domains and move beyond narrow artificial intelligence toward true AGI.
@@
  • Resoning: they often generate unreliable reasoning results, producing factually incorrect or fabricated content with apparent confidence [reasoning].
@@
  • Learning: they cannot improve or update their behavior based on new experiences or interactions, and rely entirely on retraining or fine-tuning to adjust their responses [lora].
@@
  • Memory: they cannot memorize new knowledge by themselves and must rely on external databases or retrieval systems to store and access updated information [memory].
@@
Overall, the core limitation of existing FMs is that they fail to represent the full spectrum of intelligence due to training to predict patterns in a single domain [embers]. For example, LLMs fail at learning because language alone serves as a descriptive medium rather than a functional substrate of learning. While linguistic representations can articulate the principles of learning, they cannot embody the complex neural dynamics through which biological learning actually occurs. As shown in Fig. 1, LLMs only learn statistical regularities and contextual dependencies in human language to predict and generate coherent text; time series FMs capture temporal dependencies in financial [financial] and weather data [weather]; vision FMs model spatial, temporal, and semantic continuity in images [image] and videos [video]; and world models [world] learn the dynamics and causal relationships of physical environments. In general, existing FMs primarily capture the partial expressions or surface regularities of intelligence, and thus fail to capture its underlying mechanisms comprehensively [tool, pinker].
@@
Key advantages of IFM for AGI: Unlike existing FMs that are designed for specific modalities such as language or vision, IFM learns directly from diverse human behaviors that embody broad intelligence. Intelligence, however, has long been an elusive subject, and numerous theoretical frameworks have attempted to explain its nature. For example, predictive processing theories, including Free Energy Principle (FEP) [fep] and the predictive mind hypothesis [predict], conceptualize the brain as a predictive machine that seeks to minimize the discrepancy between internal predictions and sensory observations. Meanwhile, information-integration theories, such as Global Workspace Theory (GWT) [gwt1], emphasize that intelligence emerges from the integration of specialized modules that share and broadcast information across a global neural workspace [gwt2]. Other approaches focus on more localized mechanisms, such as the complementary learning system frameworks, which highlight the distinct yet cooperative roles of the hippocampus and neocortex in forming long- and short-term memories [hippo], or studies that track how the hippocampus and striatum encode value signals during decision-making [hippo2]. While these theories offer valuable insights into specific aspects of cognition, most focus on explaining why intelligence arises rather than how to model and replicate it in machines at scale for building AGI.
@@
It is widely acknowledged that human intelligence arises from the collective dynamics of approximately 86 billion neurons in the brain [86]. However, the immense complexity of these interactions makes it nearly impossible to construct an explicit, mechanistic model of intelligence from first principles. To address this challenge, IFM is designed to capture the implicit structural neuronal activity that underlies observed intelligent behaviors, learning directly from large-scale neuronal data reflecting diverse human cognitive processes. By grounding itself in observable behaviors rather than abstract theoretical formulations, IFM avoids reliance on hand-crafted models of intelligence and offers a more practical pathway toward AGI.
@@
Conceptual Overview of IFM: As illustrated in Fig. 2, IFM learns the implicit structural neuronal activity that underlies observed intelligent behaviors. It achieves this through two core designs:
  • State Neural Network: IFM is constructed upon a state neural network designed to emulate the dynamic behavior of biological neurons. This design encompasses three key aspects. First, each state neuron maintains an internal state that stores historical information for future decision-making (Neuron Function). Second, the neurons are interconnected through recurrent, graph-like connections that allow information to branch, merge, and loop across multiple pathways, supporting a wide spectrum of intelligent functions (Neuron Connectivity). Third, connection strengths are allowed to evolve based on neuronal activity, enabling the network to self-organize, adapt, and continuously refine its internal structure through experience (Neuron Plasticity).
  • Neuron Output Prediction: The collective dynamics of neurons underlying diverse human intelligent behaviors are formalized as the process of generating neuronal outputs from temporal neuronal input signals of the biological neuron system. Following this principle, IFM formulates its learning objective as predicting biological neuronal outputs from corresponding input dynamics over time, thereby enabling the model to internalize the essential computational processes that underpin intelligence. This objective directly links low-level neuronal activity to high-level intelligent behaviors, enabling the model to learn temporal, recurrent, and context-dependent transformations that characterize human intelligence. In essence, it bridges neural dynamics with the mechanisms of intelligence.
Equipped with these designs, IFM can be trained using structured datasets that encode intelligent behaviors as temporal sequences of neuronal inputs and outputs, allowing the model to learn through standard backpropagation techniques.
@@ II IFM Technical Details
IFM learns the implicit structural neuronal activity that underlies observed intelligent behaviors. We detail its design as follows.
@@ II-A State Neural Network
With these neurons, state neural network $\mathcal{G}$ receives an input stimulus sequence with input neurons $(\mathbf{x}^{1},\mathbf{x}^{2},…,\mathbf{x}^{T}),\ \forall\mathbf{x}^{t}\in\mathbb{R}^{|\mathcal{N}_{in}|}$ over time, and generates an output response sequence through output neurons $(\mathbf{y}^{1},\mathbf{y}^{2},…,\mathbf{y}^{T}),\ \forall\mathbf{y}^{t}\in\mathbb{R}^{|\mathcal{N}_{out}|}$. By that, the function of $\mathcal{G}$ can be defined as $\mathbf{y}^{t}=\mathcal{G}(\mathbf{x}^{1},\mathbf{x}^{2},...,\mathbf{x}^{t}),\ \forall 1\leq t\leq T$, as shown in Fig. 3.
@@
The design of the state neural network $\mathcal{G}$ models the neuron dynamics of biological neural systems, and it is built around three key components: neuron function, neuron connectivity, and neuron plasticity. Neuron function captures the internal state and temporal evolution of biological neurons, allowing each neuron to integrate signals over time and retain memory of past activations. Neuron connectivity models the flexible and recurrent wiring of the brain neuron cells, enabling signal flows through complex graphs to support oscillations, synchronization, and context-dependent computation. Neuron plasticity captures the adaptive adjustment of connection strengths in response to neuronal activity, reflecting biological learning mechanisms. Together, these three components establish a computational framework that embodies the dynamic, adaptive, and self-organizing nature of intelligence. They are detailed as follows.
@@
Neuron Function: The function of a state neuron $\eta\in\mathcal{N}$ is to maintain a dynamic internal state $s$ that integrates past information and transforms incoming signals into outputs over time, thereby enabling temporal and context-dependent behavior. This process can be formalized as $v^{t}=\phi(\mathbf{u}^{t},s^{t-1})$, where $\mathbf{u}^{t}$ and $v^{t}$ denote the input and output at time $t$, respectively. In this formulation, the neuron state $s^{t}=\psi(\mathbf{u}^{t},s^{t-1})$ recursively integrates information from all previous time steps up to $t$. Recurrent [recurrent] and spiking neurons [spiking] can be regarded as two special cases within this general structure. Specifically, a recurrent neuron produces continuous outputs, with its internal state updated through a nonlinear transformation $s^{t}=\sigma(W_{\text{in}}\mathbf{u}^{t}+W_{s}s^{t-1})$, where $W_{\text{in}}$ and $W_{s}$ are the input and recurrent weight matrices, respectively. In contrast, a spiking neuron generates discrete spikes, with its internal state governed by membrane potential dynamics. For example, in the Leaky Integrate-and-Fire (LIF) model [lif], the membrane potential evolves as $\frac{ds(t)}{dt}=-(s(t)-s_{\text{rest}})+\mathbf{w}^{\top}\mathbf{u}(t)$, where $s_{\text{rest}}$ is the resting potential and $\mathbf{w}$ denotes the synaptic weights for each input.
@@
Neuron Connectivity: In the state neuron network $\mathcal{G}$, neurons are interconnected through directed, weighted edges represented by an adjacency matrix $\mathcal{E}\in\mathbb{R}^{\mathcal{N}\times\mathcal{N}}$, where $e_{ij}$ denotes the weight of the edge from neuron $\eta_{i}$ to neuron $\eta_{j}$. Unlike the layer-wise connectivity typically adopted in existing FMs, our approach employs a graph-structured state neural network. This flexible connectivity supports both feedforward pathways and recurrent loops, allowing information to branch, merge, and circulate throughout the network. Consequently, signals can interact dynamically across time and multiple pathways, enabling the model to capture temporal dependencies, integrate heterogeneous information, and coordinate complex behaviors. These are capabilities that traditional layer-wise architectures inherently struggle to achieve.
@@
Note that recurrent connectivity is essential for modeling human intelligence, yet it is largely overlooked in existing FMs. Fundamentally, recurrent connectivity allows information to persist and interact over time, creating temporal continuity that is critical for integrating past experiences, forming context-dependent representations, and coordinating complex behaviors. At the lowest level, two reciprocally connected neurons can form an oscillatory loop that generates periodic signals, similar to pacemaker neurons whose intrinsic rhythms coordinate neural processes throughout the brain [pacemaker]. When scaled up, larger neuronal assemblies sustain rhythmic firing patterns that synchronize activity across populations, enabling higher-order cognition [rhythms]. These dynamics mirror the neural oscillatory activity, such as gamma and theta brain waves (generated by excitatory–inhibitory loops in the cortex and hippocampus) [brainwave].
@@
We believe that such recurrent structures will also play a crucial role in the emergence of machine consciousness. We conceptualize machine consciousness as the experience arising from continuous neural dynamics sustained by ongoing loops of information flows within the system. Through this recurrent process, external information gathered via sensory inputs is continuously integrated, giving rise to a coherent and dynamically evolving experience. Apparently, LLM does not have such machine consciousness.
@@
Neuron Plasticity: Neuron plasticity is a fundamental mechanism in biological neurons, supporting essential capabilities such as learning and memory [hebb]. In biological neural systems, the strength and efficacy of synaptic connections can change over time in response to experience, sensory inputs, or environmental demands. This adaptability allows neural circuits to encode information, form memories, and reorganize themselves to optimize performance [ltp].
@@
For two neurons in the state neural network $\mathcal{G}$, $\eta_{p}$ and $\eta_{q}$, that are connected by an edge of weight $e_{pq}$, we define $\eta_{p}$ as the source (presynaptic) neuron that transmits a signal, and $\eta_{q}$ as the target (postsynaptic) neuron that receives it. We see the core of neuron plasticity as the edge weights are affected by the historical activities of connected neurons. That is, the weight $e_{pq}$ is influenced by the outputs of both neurons, i.e., $v_{p}$ of $\eta_{p}$ (or $u_{q}$ of $\eta_{q}$, equivalently) and $v_{q}$ of $\eta_{q}$. Then, we have $e_{pq}^{t}=\psi_{e}(u_{q}^{1},...,u_{q}^{t},v_{q}^{1},...,v_{q}^{t})$. Since $v_{q}=\phi(u_{q},s_{q})$, we integrate it and the weight update becomes $e_{pq}^{t}=\psi_{e}(u_{q}^{t},e_{pq}^{t-1})$, with $e_{pq}^{t-1}$ as the previous edge weight. This can easily extend to target neurons with multiple inputs, i.e., $e_{pq}^{t}=\psi_{e}(\mathbf{u}_{q}^{t},e_{pq}^{t-1})$.
@@
Many existing neural plasticity models fit such a formulation. For example, the widely adopted Spiking-Timing-Dependent-Plasticity (STDP) [stdp] in spiking neural networks [spike] can be fitted as $e_{pq}^{t}=\psi_{e}(\mathbf{u}_{q}^{t},e_{pq}^{t-1})=e_{pq}^{t-1}+\Delta e_{pq}^{t}$, with $\Delta e_{pq}^{t}$ determined by:
@@
  • Long-term Potentiation (LTP): If $\eta_{p}$ fires before $\eta_{q}$ (within a certain time window), $\Delta e_{pq}^{t}$ is a positive weight increase.
@@
  • Long-term Depression (LTD): If $\eta_{p}$ fires after $\eta_{q}$, $\Delta e_{pq}^{t}$ is a negative weight decrease.
@@
Summary: With the above neuron function, neuron connectivity and neuron plasticity, the model of the entire state neural network becomes $\mathbf{y}^{t}=\mathcal{G}(\mathbf{x}^{t},(\mathbf{s}^{t-1},\mathbf{e}^{t-1})),\ \forall 1\leq t\leq T$. $\mathbf{s}^{t-1}$ represents states of all neurons and $\mathbf{e}^{t-1}$ all edge weights at time t1t-1, with $\mathbf{s}^{t}=\psi(\mathbf{u}^{t},\mathbf{s}^{t-1})$ and $\mathbf{e}^{t}=\psi_{e}(\mathbf{u}^{t},\mathbf{e}^{t-1})$.

### v3

@@ Abstract
We propose a new perspective for approaching artificial general intelligence (AGI) through the Intelligence Foundation Model (IFM). Unlike existing foundation models (FMs), which learn statistical regularities within isolated manifestations of intelligence, such as language or vision, IFM aims to infer the structural principles from which these manifestations arise. This perspective is grounded in the view that intelligence emerges from stable neural architectures shaped by millions of years of evolution, enabling intelligent behavior across diverse environments. By learning from a broad spectrum of intelligent behavior rather than any single domain, IFM seeks to recover the structural principles that give rise to intelligence itself. For this perspective, IFM introduces two core components: a novel architecture, State Neural Network (SNN), and a new training objective, neuron output prediction. SNN models the states, connectivity, and plasticity of biologically inspired neurons, supporting the system to store, integrate, and process information over time. The neuron output prediction objective provides a unified computational principle for learning the structural neural dynamics that underlie diverse intelligent behavior. Together, these innovations establish a biologically grounded and computationally feasible foundation for building systems capable of generalization, reasoning, and adaptive learning across domains, making a step toward AGI.
@@ I Introduction
Artificial general intelligence (AGI) refers to systems capable of understanding, learning, and generating adaptive behavior across diverse tasks and environments, comparable to or exceeding human-level intelligence [agi]. Despite decades of progress, AGI has not yet been realized. To advance toward AGI, we propose a fundamentally different perspective through an Intelligence Foundation Model (IFM). Rather than treating intelligence as a set of task-specific abilities (e.g., language, vision, or motor skills), IFM views intelligence as an emergent property of stable neural dynamics honed by millions of years of evolution, enabling adaptive behavior in the physical world. Human brains are the best-characterized systems that instantiate such intelligence [humanagi]. While individual brains differ anatomically, they all realize a shared manifold of neural structure that supports general intelligent behavior across individuals. IFM therefore aims to learn the underlying structural principles that give rise to human intelligence. Meanwhile, any intelligent behavior, whether perception, reasoning, planning, or decision-making, can ultimately be expressed as temporal transformations from neuronal inputs to outputs. Therefore, understanding intelligence reduces to understanding the neural dynamics that give rise to these transformations. In this view, IFM reframes the challenge of “learning intelligence” as the problem of capturing the structural organization underlying intelligent neural dynamics. The shift from a task-driven paradigm to a structure-driven paradigm allows IFM to focus on the implicit principles that evolution has discovered for building general-purpose intelligent systems. Following the foundation model paradigm [fm], IFM leverages large-scale neuronal input-output transformation data that comprehensively represent diverse human behavior to support this endeavor, thereby advancing toward AGI.
@@
  • Reasoning: FMs often generate unreliable reasoning results, producing factually incorrect or fabricated content with apparent confidence [reasoning].
@@
  • Learning: FMs cannot update behavior based on new experiences and rely entirely on retraining or fine-tuning [lora].
@@
  • Memory: FMs cannot autonomously memorize new knowledge, relying on external databases or retrieval systems [memory].
@@
A core limitation of existing FMs is that they only learn patterns within specific domains, instead of capturing the full spectrum of intelligence [embers]. For example, LLMs fail at learning because language alone serves as a descriptive medium rather than a functional substrate for learning. While linguistic representations can articulate the principles of learning, they cannot embody the complex neural dynamics through which biological learning occurs. As shown in Fig. 1, LLMs capture statistical regularities in human language; time series FMs model temporal dependencies in financial [financial] or weather data [weather]; vision FMs encode spatial, temporal, and semantic continuity in images [image] and videos [video]; and world models [world] learn the dynamics and causal relationships of physical environments. Overall, existing FMs capture only partial expressions or surface regularities of intelligence, but fail to model its underlying principles comprehensively [tool, pinker].
@@
Key advantages of IFM for AGI: Unlike existing FMs, which are designed for specific modalities such as language or vision, IFM learns directly from diverse human behavior that embodies broad intelligence. However, intelligence has long been an elusive subject, and numerous theoretical frameworks have attempted to explain its nature. Predictive processing theories, including Free Energy Principle (FEP) [fep] and the predictive mind hypothesis [predict], conceptualize the brain as a predictive machine that seeks to minimize the discrepancy between internal predictions and sensory observations. Information-integration theories, such as Global Workspace Theory (GWT) [gwt1], emphasize that intelligence emerges from the integration of specialized modules that share and broadcast information across a global neural workspace [gwt2]. Other approaches focus on more localized mechanisms, such as the complementary learning systems, which highlight the distinct yet cooperative roles of the hippocampus and neocortex in forming long- and short-term memories [hippo], or studies that track how the hippocampus and striatum encode value signals during decision-making [hippo2]. While these theories offer valuable insights into aspects of cognition, they primarily explain why intelligence arises rather than how to model and replicate it in machines at scale for building AGI.
@@
Human intelligence emerges from the collective dynamics of roughly 86 billion neurons [86]. However, the immense complexity of these interactions makes it nearly impossible to construct an explicit, mechanistic model of intelligence from first principles. To address this challenge, IFM is designed to capture the implicit structural neuronal dynamics that underlie observed intelligent behavior, learning directly from large-scale neuronal data representing diverse human cognitive processes. By grounding itself in observable behavior rather than abstract theoretical formulations, IFM avoids reliance on hand-crafted models of intelligence and offers a more practical pathway toward AGI.
@@
Conceptual Overview of IFM: As illustrated in Fig. 2, IFM learns the implicit structural neuronal dynamics that underlie observed intelligent behavior. It achieves this through two core designs:
  • State Neural Network: IFM is constructed upon a state neural network designed to model three key dynamic behaviors of biological neurons. First, each state neuron maintains an internal state that encodes past activity to guide future decisions (Neuron Function). Second, the neurons are interconnected through recurrent, graph-like connections that allow information to branch, merge, and loop across multiple routes, supporting a wide spectrum of intelligent functions (Neuron Connectivity). Third, connection strengths are allowed to evolve based on neuronal dynamics, allowing the network to self-organize, adapt, and continuously refine its internal structure through experience (Neuronal Plasticity).
  • Neuron Output Prediction: The collective dynamics underlying diverse human intelligent behavior are formalized as generating neuronal outputs from the temporal input signals of biological neural systems. Building on this formulation, IFM defines its training objective as predicting biological neuronal outputs from the corresponding input neuronal dynamics over time, thereby enabling the model to internalize the essential computational processes that underpin intelligence. This objective directly links low-level neuronal dynamics to high-level intelligent behavior, enabling the model to learn temporal, recurrent, and context-dependent transformations that characterize human intelligence. In essence, it bridges neural dynamics with the principles of intelligence.
Building on these designs, IFM learns intelligence from structured datasets that encode intelligent behavior as temporal sequences of neuronal inputs and outputs, allowing training through standard backpropagation.
@@ II IFM Technical Details
IFM models the structural neuronal dynamics that underlie observed intelligent behavior. Its design is detailed as follows.
@@ II-A State Neural Network
With these neurons, state neural network $\mathcal{G}$ receives an input stimulus sequence with input neurons $(\mathbf{x}^{1},\mathbf{x}^{2},…,\mathbf{x}^{T}),\ \forall\mathbf{x}^{t}\in\mathbb{R}^{|\mathcal{N}_{in}|}$ over time, and generates an output response sequence through output neurons $(\mathbf{y}^{1},\mathbf{y}^{2},…,\mathbf{y}^{T}),\ \forall\mathbf{y}^{t}\in\mathbb{R}^{|\mathcal{N}_{out}|}$, where $T$ is the timestamp. By that, the function of $\mathcal{G}$ can be defined as $\mathbf{y}^{t}=\mathcal{G}(\mathbf{x}^{1},\mathbf{x}^{2},...,\mathbf{x}^{t}),\ \forall 1\leq t\leq T$.
@@
The design of the state neural network $\mathcal{G}$ models the neuron dynamics of biological neural systems, based on three key components: neuron function, neuron connectivity, and neuronal plasticity. Neuron function captures the internal state and temporal evolution of biological neurons, enabling each neuron to integrate inputs and maintain memory over time. Neuron connectivity models the flexible, recurrent wiring of neurons in the brain, enabling signals to propagate through complex recurrent graphs that support oscillations, synchronization, and context-dependent computation. Neuronal plasticity captures the adaptive modulation of connection strengths in response to neuronal dynamics, reflecting biological learning mechanisms. Together, these three components establish a computational framework that embodies the dynamic, adaptive, and self-organizing nature of intelligence, as shown in Fig. 3.
@@
Neuron Function: The function of a state neuron $\eta\in\mathcal{N}$ is to maintain a dynamic internal state $s$ that integrates past information and transforms incoming signals into outputs over time, thereby enabling temporal and context-dependent behavior. This process can be formalized as $v^{t}=\phi(\mathbf{u}^{t},s^{t-1})$, where $\mathbf{u}^{t}$ and $v^{t}$ denote the input and output at time $t$, respectively. In this formulation, as shown in Fig. 3, the neuron state $s^{t}=\psi_{s}(\mathbf{u}^{t},s^{t-1})$ recursively integrates information from all previous time steps up to $t$. Recurrent [recurrent] and spiking neurons [spiking] can be regarded as two special cases within this general structure. Specifically, a recurrent neuron produces continuous outputs, with its internal state updated through a nonlinear transformation $s^{t}=\sigma(W_{\text{in}}\mathbf{u}^{t}+W_{s}s^{t-1})$, where $W_{\text{in}}$ and $W_{s}$ are the input and recurrent weight matrices, respectively. In contrast, a spiking neuron generates discrete spikes, with its internal state governed by membrane potential dynamics. For example, in the Leaky Integrate-and-Fire (LIF) model [lif], the membrane potential evolves as $\frac{ds(t)}{dt}=-(s(t)-s_{\text{rest}})+\mathbf{w}^{\top}\mathbf{u}(t)$, where $s_{\text{rest}}$ is the resting potential and $\mathbf{w}$ denotes the synaptic weights for each input.
@@
Neuron Connectivity: In state neuron network $\mathcal{G}$, neurons are interconnected through directed, weighted edges represented by an adjacency matrix $\mathcal{E}\in\mathbb{R}^{\mathcal{N}\times\mathcal{N}}$, where $e_{pq}$ denotes the weight of the edge from neuron $\eta_{p}$ to neuron $\eta_{q}$. Unlike the layer-wise connectivity typically adopted in existing FMs, our approach employs a graph-structured state neural network. This flexible connectivity supports both feedforward pathways and recurrent loops, allowing information to branch, merge, and circulate throughout the network. This allows signals to interact dynamically across time and pathways, enabling the model to capture temporal dependencies, integrate heterogeneous information, and coordinate complex behavior. These are capabilities that traditional layer-wise architectures inherently struggle to achieve.
@@
Notably, recurrent connectivity is essential for modeling human intelligence but is largely overlooked in existing FMs. It enables information to persist and interact over time, creating temporal continuity that is critical for integrating past experiences, forming context-dependent representations, and coordinating complex behavior. At the microcircuit level, two reciprocally connected neurons can form an oscillatory loop generating periodic signals, akin to pacemaker neurons whose intrinsic rhythms coordinate brain-wide activity [pacemaker]. Scaling up, larger neuronal assemblies sustain rhythmic firing patterns that synchronize activity across populations to support higher-order cognition [rhythms]. Such dynamics mirror the neural oscillatory activity, such as gamma and theta brain waves (generated by excitatory–inhibitory loops in the cortex and hippocampus) [brainwave].
@@
We argue that such recurrent structures are likely to play a crucial role in the emergence of machine consciousness. We conceptualize machine consciousness as the experience emerging from continuous neural dynamics sustained by recurrent loops of information flow. Such recurrence continuously integrates external sensory information, producing a coherent, continuously evolving experience. Apparently, LLM does not have such machine consciousness.
@@
Neuronal Plasticity: Neuronal plasticity is a fundamental mechanism in biological neurons, supporting essential capabilities such as learning and memory [hebb]. In biological neural systems, the strength and efficacy of synaptic connections can change over time in response to experience, sensory inputs, or environmental demands. This adaptability allows neural circuits to encode information, form memories, and reorganize themselves to optimize performance [ltp].
@@
Considering two neurons in state neural network $\mathcal{G}$, $\eta_{p}$ and $\eta_{q}$, that are connected by an edge of weight $e_{pq}$, we define $\eta_{p}$ as the source (presynaptic) neuron that transmits a signal, and $\eta_{q}$ as the target (postsynaptic) neuron that receives it. Neuronal plasticity is realized as the modulation of edge weights based on the historical activity of connected neurons. That is, the weight $e_{pq}$ is influenced by the outputs of both neurons, i.e., $v_{p}$ of $\eta_{p}$ (or $u_{q}$ of $\eta_{q}$, equivalently) and $v_{q}$ of $\eta_{q}$. Then, we have $e_{pq}^{t}=\psi_{e}(u_{q}^{1},...,u_{q}^{t},v_{q}^{1},...,v_{q}^{t})$. Substituting $v_{q}=\phi(u_{q},s_{q})$, the update can be expressed recursively as $e_{pq}^{t}=\psi_{e}(u_{q}^{t},e_{pq}^{t-1})$, where $e_{pq}^{t-1}$ as the previous edge weight. This formulation naturally extends to target neurons with multiple inputs, i.e., $e_{pq}^{t}=\psi_{e}(\mathbf{u}_{q}^{t},e_{pq}^{t-1})$.
@@
Many existing neural plasticity models can be expressed in this framework. For example, the widely adopted Spiking-Timing-Dependent-Plasticity (STDP) [stdp] in spiking neural networks [spike] can be formulated as $e_{pq}^{t}=\psi_{e}(\mathbf{u}_{q}^{t},e_{pq}^{t-1})=e_{pq}^{t-1}+\Delta e_{pq}^{t}$, where $\Delta e_{pq}^{t}$ is given by:
@@
  • Long-term Potentiation: If $\eta_{p}$ fires shortly before $\eta_{q}$ (within a certain time window), $\Delta e_{pq}^{t}>0$.
@@
  • Long-term Depression: If $\eta_{p}$ fires shortly after $\eta_{q}$ (within a certain time window), $\Delta e_{pq}^{t}<0$.
@@
Summary: With the above neuron function, neuron connectivity, and neuronal plasticity, the model of the state neural network becomes $\mathbf{y}^{t}=\mathcal{G}(\mathbf{x}^{t},(\mathbf{s}^{t-1},\mathbf{e}^{t-1})),\ \forall 1\leq t\leq T$. $\mathbf{s}^{t-1}$ represents states of all neurons, and $\mathbf{e}^{t-1}$ denotes all edge weights at time t1t-1, with $\mathbf{s}^{t}=\psi_{s}(\mathbf{u}^{t},\mathbf{s}^{t-1})$ and $\mathbf{e}^{t}=\psi_{e}(\mathbf{u}^{t},\mathbf{e}^{t-1})$.